OpenStudy (anonymous):

One number is 15 more than another. The sum of twice the larger and 3 times the smaller is 182. find the numbers. 2 variables

OpenStudy (campbell_st):

the numbers are x and x + 15 then 3(x) + 2(x + 15) = 182 then 3x + 2x + 30 = 182 5x = 152 x = 30.4 x + 15 = 45.4 the numbers are 30.4 and 45.4

OpenStudy (anonymous):

yea but you have to use 2 variables like x and y

OpenStudy (campbell_st):

write it in 1 variable.... x the smaller and x + 15 the larger.... its easier than 3x + 2y = 182 y = x + 15 then its still the same problem using substitution 3x +2(x +15) = 182 5x + 30 = 182